# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v2

@@ 4.1 MDP formulation
The solution construction procedure by GFlowNets can be modeled as a Markov decision process (MDP). We illustrate this procedure using the Traveling Salesman Problem (TSP) as an example. This formulation can be easily adapted to general combinatorial optimization problem with minor modifications.
@@
State $s$: The initial state $s_{0}$ is an empty set. The state $s_{t}$ at $0<t<n$ is a partial combinatorial solution of prensented with previous actions: $\{a_{0},\ldots,a_{t-1}\}$ and $s_{n}$ is completed solution.
@@
Action $a$: The action $a_{t}$ is a selection of a node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$. If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x(\mathcal{G})$.
@@
Reward $R(x(\mathcal{G});\beta)$: The reward is defined as the unnormalized Boltzmann distribution of the solution $x(\mathcal{G})$ with energy $\mathcal{E}(x(\mathcal{G}))$. Specifically, the reward is given by $R(x(\mathcal{G});\beta)=e^{-\beta\mathcal{E}(x(\mathcal{G}))}$. Here, $\mathcal{E}(x(\mathcal{G}))$ represents the energy of the solution $x(\mathcal{G})$, which corresponds to metrics such as the tour length in the TSP.
The coefficient $\beta$, known as the inverse temperature, is a hyperparameter.
4.2 Policy configuration
@@
We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge matrix $\eta(\mathcal{G};\theta)$, following Ye et al. [79]. Then, the forward policy when $t<N$ is expressed as follows:
$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho)\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G};\theta)&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (5)
@@
Note that $\rho$ and $\eta(\mathcal{G};\theta)$ denote the pheromone trails and learned heuristic matrix in ACO. Subsequently, we can define a trajectory from the initial state to a terminal solution, with $\tau=(s_{0},\ldots,s_{N},x(\mathcal{G}))$, and the forward probability of that trajectory,
$P_{F}(\tau;\eta(\mathcal{G};\theta),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G};\theta),\rho).$ (6)
Note that the forward probability for the transition from $s_{N}$ to $x(\mathcal{G})$ is omitted as it follows a deterministic rule.
For the backward policy, we set $P_{B}=\frac{1}{2N}$, emulating the maximum entropy GFlowNets [83] to make uniform exploration over symmetric trajectories which terminate into identical solution. Here, $P_{B}$ represents a uniform distribution across all $2N$ feasible trajectories leading to a solution $x(\mathcal{G})$. This is particularly pertinent in the context of the TSP involving $N$ cities, where $2N$ permutations culminate in the same Hamiltonian cycle $x(\mathcal{G})$. The count of symmetric trajectories depends on the task’s characteristics. For an exhaustive discussion on solution symmetry across various COPs, please refer to Appendix E.
4.3 Solution sampling with GFACS
In the sampling process, we assume the existence of the learned GNN parameterized with $\theta^{*}$; see Section 5 for details of the training procedure.
The learned GNN provides the heuristic matrix $\eta(\mathcal{G};\theta^{*})$ for ACO algorithm. Thus, the remaining process is to update pheromone parameter $\rho$ with solutions sampled by $P_{F}(\tau;\eta(\mathcal{G};\theta^{*}),\rho)$ during the solution searching process. As the $\eta(\mathcal{G};\theta^{*})$ is trained with GFlowNets to sample diverse high-rewarded (i.e., low energy) samples, $P_{F}$ generates good candidates for updating $\rho$. The overall process is illustrated in Figure 1.
@@ 5 Off-policy training of GFACS
This section introduces how to train GFACS, especially the GNN component parameterized with $\theta$, in the GFlowNet framework. GFACS is trained via iterations of the off-policy experience collection (Step A) and training (Step B).
5.1 Step A: Off-policy experience collection with local search and energy reshaping
Figure 2: Hindsight observation with local search and energy reshaping to control the trade-off between exploration and exploitation.
@@ 5.1 Step A: Off-policy experience collection with local search and energy reshaping
In this step, we collect experiences on given graph instances. The experiences in step A will be used for training GFACS in step B. For every iteration, we first sample $M$ problem instances from a random sampler $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$. Details of instance sampling are described in Appendix B. For each problem instance $\mathcal{G}_{m}$, we collect $K$ solutions by following Eq. 6 as $\mathcal{B}_{\text{explore}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x_{m,k}},\mathcal{E}(x_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$.
@@
To exploit high-reward regions in the large-scale combinatorial space, we also collect $K$ exploitation samples for each $\mathcal{G}_{m}$ using a local search operator ($LS$), such as 2-opt for TSP, as a powerful behavioral policy. The local search operator directly refines a given solution into a locally optimized one, i.e., $x^{\prime}\leftarrow LS(x)$. We then use the backward policy $P_{B}$ to sample a trajectory $\tau_{\rightarrow x^{\prime}}$ for each $x^{\prime}$, obtaining a batch of “exploitation experiences,” $\mathcal{B}_{\text{exploit}}=\{(\mathcal{G}_{m},\tau^{m,k}_{\rightarrow x^{\prime}_{m,k}},\mathcal{E}(x^{\prime}_{m,k})\}_{m=1,\ldots,M,k=1,\ldots,K}$. Please note that we also suggest a general-purpose local search operator that partially destroys a solution and then repairs for problems that lack domain-specific local search methods. Please refer to Appendix C for details about the destroy-and-repair local saerch.
@@
While the concept of using local search for exploitation has been explored in recent GFlowNet works [39], our distinct approach involves utilizing the local search (LS) operation to facilitate hindsight observation—a forward-looking insight derived from extending the search beyond the current state. We focus on evaluating samples that were initially overlooked due to their high energy (i.e., low reward) but can be transformed to exhibit low energy (i.e., high reward) following the local search.
@@
To achieve this, we modify the energy (in a way we call energy reshaping) for the exploration around $x$ in $\mathcal{B}_{\text{explore}}$ by using the energy of the refined solution $x^{\prime}=LS(x)$ as follows:
$\mathcal{E}_{\text{reshaped}}(x)=\alpha\mathcal{E}(x^{\prime})+(1-\alpha)\mathcal{E}(x).$ (7)
5.2 Step B: Training with collected experiences with shared energy normalization
Given the collected experience batches $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, we aim to minimize the following loss function:
$\mathcal{L}(\theta)=\mathcal{L}(\theta;\mathcal{B}_{\text{explore}})+\mathcal{L}(\theta;\mathcal{B}_{\text{exploit}}),$ (8)
where ($\mbox{where }\mathcal{L}(\theta;\mathcal{B})=\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{2NP_{F}(\tau^{m,k}_{\rightarrow x_{m,k}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m})}{R(x_{m,k};\beta)}\right)^{2}.$ (9)
@@ 5.2 Step B: Training with collected experiences with shared energy normalization
The $2N$ comes from the fact that TSP has $2N$ permutations, as discussed in Section 4.2. In addition, our target sampling distribution is proportional to reward $p(x;\mathcal{G})\propto R(x(\mathcal{G}))$. However, in most COPs, including TSP, the scale of $R(x(\mathcal{G}))$ can vary significantly across instances, possibly leading to unstable training due to the high variance of logarithms of rewards. Similarly to previous works that stabilize REINFORCE training [43, 37], we normalize the energy $\mathcal{E}=-\log R$ using the sample mean over $K$ samples for the same $\mathcal{G}_{m}$, i.e.,
$-\log\tilde{R}(x_{m,k})=\tilde{\mathcal{E}}(x_{m,k})=\mathcal{E}(x_{m,k})-\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}(x_{m,k}).$ (10)
@@
We call this technique a shared energy normalization. Note that the shared energy normalization is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different from each other, especially when a heuristic local search is applied. To this end, the gradient loss is computed as follows:
$\nabla_{\theta}\mathcal{L(\theta)}\approx\nabla_{\theta}\sum_{m=1}^{2M}\sum_{i=1}^{I}\left(\log\frac{2NP_{F}(\tau^{i,m}_{\rightarrow x_{i,m}};\eta(\mathcal{G}_{m};\theta),\rho)Z_{\theta}(\mathcal{G}_{m};\theta)}{\tilde{R}(x_{i,m})}\right)^{2}.$ (11)
@@ 6 Experiments
This section presents experimental results to validate our algorithm [^1] Source code is available: https://github.com/ai4co/gfacs. . We compare GFACS with other ACO algorithms, and with competitive reinforcement learning algorithms for vehicle routing problems. Moreover, we conduct an ablation study, and perform an empirical comparison with previous GFlowNets training methods, which serve a similar purpose to our method. Below, we provide a brief overview of the experimental setup.
@@
Implementation and hyperparameters. GFACS is built upon the implementation established by DeepACO [79]. As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z_{\theta}$ in Equation 9. For detailed network architecture, hyperparameters and hardware settings, please refer to Appendix A. It is worth noting that one important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, which enables us to control the diversity-optimality trade-offs (see Section D.2 for detailed analysis). In addition, we employ Ant System [14] for the pheromone updates throughout the experiments, while we also provide experimental results with other ACO variants in Section D.3.
@@
Test settings. Each algorithm is tested on the held-out datasets, each with 100 instances generated in advance except TSP, where we obtain the TSP test dataset with 128 instances from Ye et al. [79]. Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct random seed.
6.1 Comparison with ACO algorithms
Here, we compare GFACS against the vanilla ACO with and the DeepACO [79] algorithm across various CO tasks.
Problems. Following the approach of Ye et al. [79], we evaluate our algorithm in a range of COPs. This includes four routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), CVRP with Time Window (CVRPTW), Prize-Collecting Vehicle Routing Problem (PCTSP), and Orienteering Problem (OP) – along with a scheduling problem – Single Machine Total Weighted Tardiness Problem (SMTWTP) – and a grouping problem, the Bin Packing Problem (BPP). For detailed problem configurations, we direct readers to Appendix B. For TSP, CVRP and CVRPTW, we use local search algorithms with perturbation similar to Ye et al. [79]. For others, we use the destroy-and-repair method (see Appendix C for more details).
Figure 3: Results of ACO algorithms on various CO tasks. Our GFACS outperforms every ACO baseline. The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models.
Baselines. We use vanilla ACO with a human-designed heuristic matrix, DeepACO as baseline algorithms. We also test GFACS without the proposed off-policy training scheme (called On-policy TB). For inference, we use $K=100$ ants and $T=100$ ACO iterations for all the algorithms.
Results. Figure 3 show GFACS consistently outperforms both the classic ACO algorithm and DeepACO across all benchmark problems. Notably, for routing problems, GFACS surpasses DeepACO by a significant margin. This is particularly remarkable given that both algorithms use identical neural networks and local search operators, underscoring the importance of our training strategy. Additionally, the proposed off-policy training method enhances GFACS performance, demonstrating its effectiveness for GFlowNets.
6.2 Comparison with other reinforcement learning methods for combinatorial optimization
In this section, we present experimental results comparing our method with reinforcement learning (RL)-based approaches, as both aim to develop combinatorial solvers without the need for labeled optimal data points, unlike supervised learning methods (e.g., [16, 70]).
Baselines. For TSP baselines, we select the AM [41], POMO [43], Sym-NCO [37] and Pointerformer [28] as RL-based constructive methods, DIMES [62] as RL-based heatmap-based methods, the Select and Optimize (SO) [9] as RL-based improvement methods. We also applied two search algorithms, the Efficient Active Search (EAS) [23] and Simulation Guided Beam Search (SGBS) [10], to POMO for a more thorough comparison. For CVRP, [^2] We also adopt TAM as a CVRP baseline; however, due to the inaccessibility of source code, we compare it using the reported result in the paper [24]. See Section D.5. we employ POMO and Sym-NCO as baselines with EAS and SGBS. Note that we restrict the duration of these searches to no more than 5 and 10 minutes for problems with 500 and 1000 nodes, respectively. We use Concorde [2] for TSP and PyVRP [77] for CVRP as oracle heuristic solvers to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap. For ACO algorithms, we use 100 ants with $T=10$ with local search algorithms similar to Ye et al. [79].
Results. Among the baseline methods evaluated, our GFACS algorithm consistently delivers either superior or highly competitive performances. When compared to faster methods like AM or POMO, GFACS demonstrates substantially better results. Even after an additional search with EAS or SGBS, they fail to match GFACS’s performance. Compared to algorithms like DIMES and Pointformer, which operate at similar speeds, GFACS achieves improved performance. GFACS performs slightly worse than the oracle heuristic solver LKH-3, but offers an advantage of faster inference, especially in CVRP. Therefore, we can conclude that GFACS, as a general-purpose algorithm, holds significant appeal in the highly competitive benchmarks of both TSP and CVRP.
@@ 6.2 Comparison with other reinforcement learning methods for combinatorial optimization
Table 1: Experimental results on TSP and CVRP. Results for methods with * are drawn from Ye et al. [79], Cheng et al. [8], and Jin et al. [28]. The reported objective value is averaged over 128 instances of TSP and 100 instances of CVRP. We report the average computation time per instance.
@@
Table 2: Experimental results on real-world datasets, TSPLib and CVRPLib-X. ‘#’ denotes the number of instances in each set. We report the average optimality gap using the best-known costs.
TSPLib # ACO DeepACO Ours CVRPLib-X # ACO DeepACO Ours
100-299 30 1.71% 1.25% 1.21% 100-299 43 2.50% 3.27% 2.69%
300-699 10 4.26% 2.70% 2.60% 300-699 40 3.75% 4.22% 3.71%
700-1499 12 7.01% 4.08% 3.88% 700-1001 17 4.64% 5.07% 4.32%

### v4

@@ 3.1 MDP formulation
We illustrate the solution construction procedure as a Markov decision process (MDP), using the Traveling Salesman Problem (TSP) as an example. This formulation can be easily adapted to general combinatorial optimization problems with minor modifications.
@@
State $s$: The initial state $s_{0}$ is an empty set. The state $s_{t}$ at $0<t<n$ is a partial combinatorial solution presented with previous actions: $\{a_{0},\ldots,a_{t-1}\}$ and $s_{n}$ is a completed solution.
@@
Action $a$: The action $a_{t}$ is a selection of the next node to visit at time $t$ among un-visited nodes: $a_{t}\in\{1,\ldots,N\}\setminus\{a_{0},\ldots,a_{t-1}\}$. If all nodes are visited, a deterministic terminating action is applied, which maps a final state $s_{N}$ into a solution, $x$.
@@
Reward $R$: The reward is given by $R(x;\mathcal{G},\beta)=e^{-\beta\mathcal{E}(x;\mathcal{G})}$. Here, $\mathcal{E}(x;\mathcal{G})$ represents the energy of the solution $x$, which corresponds to metrics such as the tour length in the TSP. The inverse temperature $\beta$ is a hyperparameter. For simplicity, we will omit $\mathcal{G}$ from $R$ and $\mathcal{E}$ except when its inclusion is necessary for clarity.
3.2 Policy configuration for prior modeling
@@
We parameterize the forward policy $P_{F}$ using a graph neural network (GNN) with parameter $\theta$ that maps an input graph instance $\mathcal{G}$ to the output edge matrix $\eta(\mathcal{G};\theta)$, following Ye et al., (2023). For simplicity, we use notation $\eta(\mathcal{G})$ without $\theta$. With a given $\eta(\mathcal{G})$, the forward policy when $t<N$ is expressed as follows:
$P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G}),\rho)\propto\begin{cases}\rho_{a_{t-1},a_{t}}\eta_{a_{t-1},a_{t}}(\mathcal{G})&\text{if $a_{t}\notin s_{t}$ }\\ 0&\text{otherwise}\end{cases}$ (5)
@@
Note that $\rho$ and $\eta(\mathcal{G})$ denote the pheromone trails and learned heuristic matrix in ACO. Subsequently, we can define a trajectory from the initial state to a terminal solution, with $\tau=(s_{0},\ldots,s_{N},x)$, and the forward probability of that trajectory as
$P_{F}(\tau;\eta(\mathcal{G}),\rho)=\prod_{t=0}^{N-1}P_{F}(s_{t+1}|s_{t};\eta(\mathcal{G}),\rho).$ (6)
The forward probability for the transition from $s_{N}$ to $x$ is omitted since it is deterministic. During training, we fix $\rho$ to $\mathbf{1}$, and thus we omit it from subsequent sections unless its presence is required for clarity.
For the backward policy, we first define $h(x)$ as the number of possible trajectories that yield the solution $x$, and we set $P_{B}(\tau|x)=\frac{1}{h(x)}$. This setting emulates the maximum entropy GFlowNets (Zhang et al., 2022a, ), treating all symmetric trajectories that terminate in the same solution equally. In the case of the TSP, where $h(x)=2N$, we have $P_{B}=\frac{1}{2N}$, thereby imposing a uniform backward distribution across all $2N$ feasible trajectories leading to the same Hamiltonian cycle $x$. It is important to note that the number of symmetric trajectories varies depending on the specific CO problem; for a more detailed discussion on solution symmetry across various CO problems, please refer to Appendix E
3.3 ACO update on learned prior
In the sampling process, we assume the existence of the learned GNN parameterized with $\theta^{*}$; see Section 4 for details of the training procedure. This learned GNN provides the prior parameter $\eta^{*}(\mathcal{G})$ for the ACO algorithm. The next step is to update the pheromone parameter $\rho$, using $K$ agents (ants) to search over samples drawn from the amortized prior $P_{F}(\tau;\eta^{*}(\mathcal{G}))$ as described in Section 2.3.
Since the trained $\eta^{*}(\mathcal{G})$ samples diverse, high-reward (i.e., low-energy) candidates, the distribution $P_{F}$ provides strong prior candidates for the ants in the ACO algorithm. This learned prior is expected to highly improve the likelihood of discovering the global optimum after the iterations of the pheromone update.
The overall process is illustrated in Figure 2.
@@ 4 Off-policy training of GFACS
This section introduces how to train GFACS, especially the GNN component parameterized with $\theta$, in the GFlowNet framework. GFACS is trained via iterations of the off-policy experience collection (Section 4.1) and training (Section 4.2).
4.1 Off-policy experience collection
Figure 3: Experience collection procedure. Energy reshaping compensates for the energy of underrated samples that have potential to become low-energy samples after the local search.
@@ 4.1 Off-policy experience collection
In this step, we collect experiences on given graph instances. The collected experiences will be used to train GFACS, as described in Section 4.2. At each training round, we first sample $M$ problem instances from a random instance generator $P(\mathcal{G})$, i.e., $\mathcal{G}_{1},\ldots,\mathcal{G}_{M}\sim P(\mathcal{G})$. Details of instance generation are described in Appendix B. For each problem instance $\mathcal{G}_{m}$, we collect $K$ solutions by following Eq.(6), resulting in a batch of “exploration experiences,”
$\mathcal{B}_{\text{explore}}=\left\{\left(\mathcal{G}_{m},\tau_{m,k},\mathcal{E}(x_{m,k})\right)\mid m\in[M],k\in[K]\right\}.$
@@
To exploit high-reward regions in the large-scale combinatorial space, we collect additional $K$ experiences for each $\mathcal{G}_{m}$ using a local search (LS) operator, such as 2-opt for TSP, as a powerful behavior policy. The LS operator directly refines a given solution into a locally optimized one, i.e., $\hat{x}\leftarrow\text{LS}(x)$. We then use the backward policy $P_{B}$ to sample a trajectory $\hat{\tau}$ for each $\hat{x}$, obtaining a batch of “exploitation experiences,”
$\mathcal{B}_{\text{exploit}}=\left\{\left(\mathcal{G}_{m},\hat{\tau}_{m,k},\mathcal{E}(\hat{x}_{m,k})\right)\mid m\in[M],k\in[K]\right\}.$
For problems that lack domain-specific LS operators, we also suggest a general-purpose destroy-and-repair method that partially destroys a solution and then repairs; see Appendix C for more details.
@@
While the concept of using local search for exploitation has been explored in recent GFlowNet works (Kim et al., 2024b, ), our distinct approach involves utilizing the local search operation to facilitate hindsight observation—a forward-looking insight derived from extending the search beyond the current state. Methodologically, we compensate for the solutions that were initially overlooked due to high energy (i.e., low reward) but that can be readily transformed into low energy (i.e., high reward) solutions through local search.
@@
To achieve this, we modify the energy (in a way we call energy reshaping) of $x$ in the exploration experiences, $\mathcal{B}_{\text{explore}}$. Specifically, for a given solution $x\in\mathcal{B}_{\text{explore}}$ and its corresponding refined solution $\hat{x}\leftarrow\text{LS}(x)$, we adjust the energy of $x$ as follows:
$\mathcal{E}^{\prime}(x)=\alpha\mathcal{E}(\hat{x})+(1-\alpha)\mathcal{E}(x).$ (7)
This reshaping captures the potential for samples to become low-energy solutions after local search.
The overall procedure of experience collection via sampling and local search, along with the energy reshaping, is illustrated in Figure 3.
4.2 Training with collected experiences
Given a batch of experiences $\mathcal{B}$, the trajectory balance (TB) loss in Eq.(1) can be rewritten as
$L(\mathcal{B})=\frac{1}{MK}\sum_{m=1}^{M}\sum_{k=1}^{K}\left(\log\frac{P_{F}(\tau_{m,k};\eta(\mathcal{G}_{m}))Z(\mathcal{G}_{m})}{(1/h(x_{m,k}))R(x_{m,k};\mathcal{G}_{m},\beta)}\right)^{2},$ (8)
where $1/h(x_{m,k})$ is to the uniform backward policy as described in Section 3.2. Since the loss is computed over multiple instances $\mathcal{G}_{m}$, the objective corresponds to the conditional trajectory balance.
Figure 4: Results of ACO algorithms with different priors on various CO tasks. Our GFACS outperforms every ACO baseline. The results are averaged over 3 independent models evaluated on the held-out test sets, and the shade indicates the min-max range of the 3 models.
@@ 4.2 Training with collected experiences
As discussed in Section 2.2, our target sampling distribution is proportional to reward, i.e., $p(x;\mathcal{G})\propto R(x;\mathcal{G},\beta)$. However, in most CO problems, including TSP, the scale of $R(x;\mathcal{G},\beta)$ can vary significantly across instances, possibly leading to unstable training due to the high variance of logarithms of rewards. Similar to previous works that stabilize REINFORCE training (Kwon et al.,, 2020; Kim et al.,, 2022), we normalize the energy $\mathcal{E}$ using the sample mean over $K$ samples for the same $\mathcal{G}_{m}$ as follows:
~$\tilde{\mathcal{E}}(x_{m,k};\mathcal{G}_{m})=\mathcal{E}(x_{m,k};\mathcal{G}_{m})-\frac{1}{K}\sum_{k=1}^{K}\mathcal{E}(x_{m,k};\mathcal{G}_{m}).$ (9)
@@
Accordingly, we have $\tilde{R}(x_{m,k};\mathcal{G}_{m},\beta)=e^{-\beta\tilde{\mathcal{E}}(x_{m,k};\mathcal{G}_{m})}$. We call this technique a shared energy normalization. Note that this technique is applied independently to the two batches, $\mathcal{B}_{\text{explore}}$ and $\mathcal{B}_{\text{exploit}}$, as the energy distribution between the two batches is often significantly different from each other, especially when a heuristic local search is applied. By applying shared energy normalization to both batches, we obtain ~$\tilde{\mathcal{B}}_{\text{explore}}$ and ~$\tilde{\mathcal{B}}_{\text{exploit}}$, in which every $\mathcal{E}$ is replaced by $\tilde{\mathcal{E}}$.
The final loss to minimize is then defined as
$\mathcal{L}(\theta):=\frac{1}{2}L(\mathcal{B}_{\text{explore}})+\frac{1}{2}L(\mathcal{B}_{\text{exploit}}),$ (10)
where $L$ is as defined in Eq.(8). Note the parameters $\theta$ govern the functions $Z(\mathcal{G}_{m})$ and $P_{F}(\cdot;\eta(\mathcal{G}_{m}))$, although this dependence has been omitted for simplicity. The gradient of the loss with respect to $\theta$ is given by $\nabla_{\theta}\mathcal{L}(\theta)$, and the parameters are optimized using stochastic gradient descent methods.
@@ 5 Experiments
This section presents experimental results to validate our algorithm. We compare GFACS against other ACO algorithms with different priors and also against competitive reinforcement learning algorithms for vehicle routing problems. Moreover, we conducted an ablation study and an empirical comparison with previous GFlowNets training methods. Last but not least, we validate the efficacy of multi-modal prior with another posterior update method beyond ACO, the active search (AS). Below, we provide a brief overview of the experimental setup.
@@
Implementation. GFACS is built upon the implementation established by DeepACO (Ye et al.,, 2023). As our methodology mainly focuses on improving training algorithms rather than network architecture, we simply adopt the GNN architecture proposed in DeepACO, except the additional parameters for partition function $Z(\cdot;\theta)$ in Eq.(8). For detailed network architecture, hyperparameters, and hardware settings, please refer to Appendix A. It is worth noting that one important hyperparameter that GFACS newly introduces is the inverse energy temperature $\beta$, which enables us to control the diversity-optimality trade-offs (see Section D.2 for detailed analysis). In addition, we employ Ant System (Section 2.3) for the pheromone updates throughout the experiments. We provide experimental results with other ACO variants in Section D.3.
@@
Test settings. Each algorithm is tested on the held-out datasets, each with 100 instances generated in advance except TSP, where we obtain the TSP test dataset with 128 instances from Ye et al., (2023). Unless otherwise stated, the results presented throughout this section are the averaged values from three independent models, each trained with a distinct seed.
Table 1: Experimental results on real-world datasets, TSPLib and CVRPLib. ‘$N$’ denotes the range of instance size and ‘#’ denotes the number of instances in each set. We report the average optimality gap using the best-known costs.
$N$ ACO DeepACO GFACS (ours)
TSPLib 100-299 30 1.71% 1.25% 1.21%
300-699 10 4.26% 2.70% 2.60%
700-1499 12 7.01% 4.08% 3.88%
CVRPLib 100-299 43 2.50% 3.27% 2.69%
300-699 40 3.75% 4.22% 3.71%
700-1001 17 4.64% 5.07% 4.32%
@@
Table 2: Experimental results on TSP and CVRP. Results for methods with * are drawn from Ye et al., (2023), Cheng et al., (2023), and Jin et al., (2023). The reported objective value is averaged over 128 instances of TSP and 100 instances of CVRP. We report the average computation time per instance.
@@
5.1 Comprison with different priors
Baselines. We compare GFACS against the vanilla ACO (heuristic prior) and the DeepACO (Ye et al.,, 2023) (on-policy RL prior) and ACO with TB prior as baseline methods. We aim to show our multi-modal prior from our off-policy GFlowNets training methods is beneficial when this is integrated with the iterative pheromone update of ACO. We use $K=100$ ants and $T=100$ ACO iterations for all algorithms.
Problems. We evaluate our algorithm in a range of CO problems, including five routing problems – TSP, Capacitated Vehicle Routing Problem (CVRP), CVRP with Time Window (CVRPTW), Prize-Collecting TSP (PCTSP), and Orienteering Problem (OP) –, a scheduling problem, the Single Machine Total Weighted Tardiness Problem (SMTWTP), and a grouping problem, the Bin Packing Problem (BPP). Detailed descriptions and configurations for each problem can be found in Appendix B. For TSP, CVRP and CVRPTW, we use local search heuristics (2-opt, Swap* and LocalSearch operator in PyVRP (Wouda et al.,, 2024), respectively) with perturbation scheme similar to Ye et al., (2023) for training GFACS and testing all the ACO algorithms. For others, we use the destroy-and-repair method. In Appendix C, further details are provided.
Results. Figure 4 shows GFACS consistently outperforms both the classic ACO algorithm and DeepACO across all benchmark problems. Notably, for routing problems, GFACS surpasses DeepACO by a significant margin. This is particularly remarkable given that both algorithms use identical neural networks and local search operators, underscoring the importance of the multi-modal prior. Additionally, the proposed off-policy training method enhances GFACS performance, demonstrating its effectiveness for GFlowNets.